OpenStudy (anonymous):

matrices, concurrent lines and unique solutions Help with distinguishing them apart

OpenStudy (anonymous):

When given this equation: \[\left[\begin{matrix}2 & -1 &0\\ 2 & 1 & 4 \\ 3 & 0 & 2 \end{matrix}\right] \left(\begin{matrix}x \\ y\\z\end{matrix}\right)=\left(\begin{matrix}0 \\ 0 \\0\end{matrix}\right)\] The relationship of |M| not equal 0 results in a unique solution. Now this matrix has a determinant 12 I think, anyway non zero, so should there be a unique solution? answer book says no, x=y=z=0 seems unique to me? No for concurrent lines to have a solution their determinant = 0 this seems confusing! Any help much appreciated

OpenStudy (anonymous):

anybody?